OpenStudy (anonymous):

Find the value of g(-2) g(-2) = 2 (-2)^2 + (-2) - 12

OpenStudy (usukidoll):

2(4)-2-12 8-2-12 = 6-12 = -6 ?

OpenStudy (anonymous):

Thank you. :D

OpenStudy (anonymous):

I have another question which is confusing: g(1 + a) = 2(1+a)^2 + )1+a) - 12

OpenStudy (usukidoll):

that should be right ... just expand the (1+a)^2 and combine like terms

OpenStudy (usukidoll):

simplify afterwards

OpenStudy (skullpatrol):

$$\Huge g(1 + a) = 2(1+a)^2 + (1+a) - 12$$

OpenStudy (usukidoll):

OH THANK GAWD THE DRAW IS BACK!

OpenStudy (usukidoll):